### v2

# A Multi-Scale Cognitive Interaction Model of Instrument Operations at the Linac Coherent Light Source

Abstract
We describe a novel multi-agent, multi-scale computational cognitive interaction model of instrument operations at the Linac Coherent Light Source (LCLS). A leading scientific user facility, LCLS is the world’s first hard x-ray free electron laser, operated by the SLAC National Accelerator Laboratory for the U.S. Department of Energy. As the world’s first x-ray free electron laser, LCLS is in high demand and heavily oversubscribed. Our overall project employs cognitive engineering methodologies to improve experimental efficiency and scientific productivity by refining experimental interfaces and workflows, simplifying tasks, reducing errors, and improving operator safety and stress levels. Our model simulates aspects of human cognition at multiple cognitive and temporal scales, ranging from seconds to hours, and among agents playing multiple roles, including instrument operator, real time data analyst, and experiment manager. The model can predict impacts stemming from proposed changes to operational interfaces and workflows. Because the model code is open source, and supplemental videos go into detail on all aspects of the model and results, this approach could be applied to other experimental apparatus and processes. Example results demonstrate the model’s potential in guiding modifications to improve operational efficiency and scientific output. We discuss the implications of our findings for cognitive engineering in complex experimental settings and outline future directions for research.
1 Introduction
LCLS, the Linac Coherent Light Source, located at the SLAC National Accelerator Laboratory, is a world-leading x-ray source, conducting hundreds of high energy x-ray experiments annually. These experiments probe underlying nano-scale mechanisms in such diverse fields as biology, chemistry, physics, materials science, and geology.[lclsoverview]. Beam time at LCLS is a unique and limited resource that must be efficiently utilized to respond to scientific needs and to maximize the impact of these ground-breaking experiments. The goal of the present effort is to apply cognitive engineering methods to optimize LCLS experimental interfaces and workflows, thereby improving the efficiency of operations and increasing the rate of production of scientific knowledge at LCLS.
Cognitive Engineering [WilsonEtAl2013] is a methodology tailored to analyze and improve settings where humans interact with engineered systems and devices, commonly referred to as “human-in-the-loop” settings. The method combines qualitative and quantitative experiments, mathematical analysis, and computer simulations, with the goal of improving the setting’s overall efficiency by making common tasks simpler and less error prone, reducing overall errors, improving human participants’ safety, and reducing their stress.
LCLS, and other scientific user facilities present a unique challenge for cognitive engineering. scientific user facilities host an ever-changing panoply of experiments, nearly all of which are unique, often actually defining the state-of-the-art. Each year LCLS hosts approximately 300 teams from around the world.
Here we briefly describe the LCLS setting and experiment workflow, and a multi-scale, multi-agent computational model that can make rough predictions about the impact of proposed changes to LCLS interfaces and workflows. The details of the model and example results are provided in three supplemental videos[vids].
Modeling of this sort is important because, due to its complexity and cost and the limited beam time, it is not feasible to run typical human-factors experiments at LCLS, except in very simple cases. A unique feature of our model is that it integrates simulations at multiple scales, and models aspects of individual cognition, teamwork, planning, and decision-making both in-the-moment and across multiple extended data-taking runs. Although the model only represents these in rough detail, it is able to make sensible predictions regarding the impact of a wide range of proposed changes. Because of the costliness and rarity of beam time, even rough predictions can be useful in increasing the efficiency of knowledge production at LCLS.
Below we focus on a description of the LCLS setting at SLAC. We then describe the model at a high level, and explain its core concepts. Details of the model, including demos of its operation and a deep dive into its code, are left to the supplemental videos [vids]. The model itself is open source and available on GitHub [githubmodel]. We conclude with thoughts on future directions and opportunities for this research approach.
2 The Setting: Machines and Instruments
LCLS is a huge system involving an enormous number of physical and software components, and multiple teams of scientists and engineers. The physical system consists of a 2 kilometer-long linear accelerator (hereafter the “accelerator” or “linac”), and a 200 meter long magnetic undulator that, taken together, form an x-ray free electron laser (XFEL). Together, the linac and undulator are often referred to as “the machine”, and produce 120 very short (50 femtosecond) and extremely bright x-ray pulses every second. This series of x-ray pulses is referred to as the “x-ray beam” or just the “beam”. After propagating several hundred meters through various diagnostics and beam shaping optics, these x-ray pulses arrive at a number of “instruments” whose purpose is to collect data about what happens when the x-ray beam interacts with materials of interest, generally called “targets” or “samples”. [^1] Although we use these words essentially synonymously, as do the LCLS users and staff, the sample, or samples, are the set of materials or items intended to be targeted in a given experiment, but at a given moment there is only one target, although it is usually part of the sample. But maintaining this distinction is unnecessarily complex for the present usages, and, indeed, for most usages. The instruments are located in “hutches”, which range in size from about a one-car to a four-car garage. [^2] Traweek’s classic ethnography of SLAC[Traweek88] provides an excellent, if somewhat outdated, depiction of the linac and of the scientists and engineers that built and operate it (esp. pp. 37-39). The same book also describes the equivalent of what we are here calling “instruments”, which are, in that book, called “detectors” (pp. 48-51 and 71-72. There is a technical difference between what we are calling an “instrument” and what, at the time of Traweek’s work were called “detectors”, but this distinction is more subtle than is worth going into here. The XFEL and LCLS did not exist at the time of Traweek’s study.
Due to the large distances, legacy control systems and differing staffing requirements, the Machine and Instruments are operated from separate control rooms. The control room operating The Machine (the linac+XFEL) is called the Accelerator Control Room (ACR). Each instrument has its own Instrument Control Room (ICR, Figure 1), located near its hutch. In the present paper we are mostly concerned with activities in the local control room of a specific instrument, called CXI, which is used for x-ray crystallography.[xcipage]
Figure 1: A typical view of an Instrument Control Room. (Photo by the authors.)
Figure 2 abstractly depicts the layout of subsystems involved in SLAC/LCLS operations, and their rough interaction. The next section provides more detail about the systems depicted in Figure 2 (see also supplemental videos [vids], [ritter2022]).
Figure 2: The Machine (linac + XFEL) delivers x-rays to an Instrument, where the x-rays strike a target producing data which is processed by an analytic pipeline, as described in the text and supplemental videos. The system is operated through a close and continuous collaboration between scientists and engineers in the ACR and ICR who play various roles (as described in the text).
3 Experiments and Measurements
An “experiment” here means the contiguous period, usually around 3-5 days, during which users are actively using an LCLS instrument to collect data. An experiment has many sub-steps. The sub-steps that we will be primarily concerned with are the contiguous periods of time, called “measurements”, during which the x-ray beam is striking a given target—or at least intended to be doing so—with the intention of collecting data about that target under given conditions. [^3] Because the focused x-ray beam typically modifies the sample almost immediately, a given measurement typically utilizes a series of putatively identical targets, for example on a tape or embedded in a stream of liquid. We do not need to distinguish these cases, and so just use the term “target” or “sample” for either of these cases, unless it is necessary to distinguish them.
Measurements can range in length from just a few seconds (if something is being briefly tested, or if something goes wrong), to 15 minutes, or sometimes much longer, as multiple samples are cycled into the path of the x-ray beam. The specific length of time that measurements take will be of central concern in our model as time is the most limiting resource at LCLS.
There are usually many dozens — perhaps hundreds — of measurements made over the several-day period of an experiment that generate the information sought by the scientists running the experiment. There are, of course, many other parts of a start-to-finish experiment, but we are only concerned at present with the few days that the team is actually at SLAC, and only when they are actively using the instrument to make measurements. Measurements generally make up the bulk of the time of an experiment.
Importantly, most experiments utilize a variety of targets, including “dummy” targets used to align and calibrate the instrument, “control” targets that have known properties, and “critical” targets for which the experiment is designed to gather new data. Because the critical targets are often costly and of limited availability, the experiment must be planned to maximize the high quality data — that is, data with high signal-to-noise ratio (SNR) — collected on these critical targets. Therefore, although an enormous amount of careful planning goes into experiment design well before the experimental team arrives at SLAC, unplanned issues always arise, so the team engages in continuous, on-site, real-time decision-making to maximize the collection of high quality data on the critical targets.
The primary decisions to be made are how long a given measurement should continue, and in what order measurements should be taken from different targets. Because beam time is limited, it seems sensible to start gathering data on critical targets as soon as possible. However, because unexpected issues often arise during the experiment’s actual beam time, it may be more sensible to use dummy or control targets until it is clear that the instrument is working well (i.e., producing high SNR data), especially if the availability of critical targets is limited. On the other hand, as the experiment nears the end of its allocated beam time, the team may wish to get as much data as possible on critical targets, even if the instrument is not producing the expected level of SNR, and especially if things did not go as planned — for example due to unexpected linac down time — one may want to get as much data as possible on critical targets in the remaining time, regardless of the quality of the data being non-optimal. This sort of tradeoff is very important, and, as will be seen in more detail later, defines the top layer of our model.
4 Experimental Teams
The scientific productivity of this multi-billion dollar system depends on a highly complex, high performance, human-machine choreography performed by combinations of internal and external scientists and engineers who, taken together, we call “experimental teams” [^4] Traweek also discusses teamwork in the SLAC context, likening scientific teams (based in the United States) to sports teams (p. 149).[Traweek88]. Teams often have a wide diversity of backgrounds (biologists, chemists, physicists, engineers, etc.) and some diversity of experience (from undergrads through professional scientists), and usually include both visiting scientists (i.e., non-SLAC/LCLS staff), and on-site SLAC/LCLS staff, including the “instrument scientists” who develop and usually operate the instruments, and the staff operators of the linac and XFEL who work in the ACR. [^5] The LCLS instrument scientists and ACR operating team do not change with each experiment, although external scientists always work with the instrument scientist in the ICRs, and often will work as well with the machine staff in the ACR on particularly complex experiments. Unless otherwise specified, we use “team” to mean the part of the team engaged in directly using and operating the CXI instrument during a particular experiment, usually operating out of the instrument control room (or remotely interacting with the team members in the CXI ICR).
LCLS hosts hundreds of experiments each year, often several different experiments each week. Although there is extensive planning for each experiment, the team involved in a given experiment really only comes physically together at LCLS for the few days during which measurements are taken. Although subsets of a given team may have long history with one another and with LCLS, the whole team has usually not previously worked together in the LCLS setting. Moreover, the team members almost always have different amounts of experience with LCLS — ranging from students who are coming to SLAC for the first time, to full time LCLS instrument scientists who maintain the instruments and are usually the ones who carry out direct control manipulations. [^6] There are often fully internal experiments where every member of the team is a full time LCLS scientist. We are not calling these out as different from typical “user” experiments, although they do differ in that the teams involved in such internal experiments are more familiar with the setting and with working together. There are often many scientists and engineers in the ICR at a given moment. We have observed experiments with as few as two, and as many as ten or more people in an ICR. There are also, commonly, remote participants, interacting by phone or via online communications such as Zoom, or through “virtual machine” screen sharing connections.
4.1 Experiment Workflow and Virtual Roles
The goal of most experiments is to get data about a particular target’s interaction with the x-ray beam, and the team is generally allocated a few days (say 24-96 hrs) of live beam time to accomplish this. As one can well imagine, there is a great deal going on during that time. For the moment, as our model is relatively rough, a general picture of what transpires in a CXI experiment will suffice (and see supplemental videos [vids]).
Although every LCLS experiment is dependent upon the x-ray beam created in The Machine, which is operated from the ACR, there comes a point where the beam is “handed off” from The Machine to The Instrument, in the specific sense that the photons are manipulated by components of the system controlled from the ICR, and no longer from the ACR. Regardless, the success of a given measurement (and so, of the whole experiment) is dependent upon creation of x-ray pulses with the desired characteristics, and then using that beam in the instrument. Therefore, a constant and complex interaction takes place between the instrument operators in the ICR, and the machine operators in the ACR in order to get the x-ray beam delivered as desired in terms of energy, pulse duration, and the many other parameters that describe the x-ray beam, and which are in constant flux. As the ACR and the ICR are about a 15 minute walk, or a 3 minute drive, from one another, this interaction takes place mostly by phone or through screen sharing. In a very real sense, the users at the instrument end of the system are at the mercy of the skill of the upstream machine operators’ ability to keep the beam coming with the correct set of parameters, so that the scientists at the instrument end can get the data they need in the allotted period. Wasted time, for whatever reason, is data lost, and given the cost and difficulty of getting beam time, losing data can mean millions of dollars, and maybe the difference between a student graduating on time or not at all. [^7] Also, of course, they are at the mercy of unanticipated problems. Much of the skill of the operators at both ends of the system is in how to deal with such problems.
4.2 Operator, Analyst, and Experiment Manager
Given the picture painted above, it should be clear that the success of an experiment rides on the team’s ability to rapidly develop an effective “team workflow” in order to conduct as many “good” measurements as possible — i.e., ones with high signal-to-noise ratios (SNR) on important targets. We identify three “virtual roles” (depicted in Figure 2) which are always represented, regardless of the number of separate people involved in an experiment: The Instrument Operator, the Data Analyst, and the Experiment Manager.
The Instrument Operator actively controls the instrument. They are usually an individual, and they are almost always seated at the set of screens that display the status of, and permit control of the hundreds of instrument settings, such as motor positions, valve and shutter settings, and so on (Figure 3). When there are multiple operators, there is almost always a primary operator. The others may be assistants, observers, apprentices, or students. The operator relies on several sorts of information in making control decisions: “Direct Control Data” is information that arises from “Primary Data Processing”, as well as information coming from the other team members and from the ACR. These are explained in more detail below.
Figure 3: A typical control window in an LCLS instrument. Most of the depicted controls are motor controls like those in our model. In addition to using the left and right buttons, the motor position can be changed by clicking in the text box containing the number, and typing in a new value. Also, for each such control, there is almost always a command-line equivalent. Which control (user interfaces (UIs) like this one, or command lines) is used for a particular purpose in a particular setting depends on many factors in the setting and the preferences of the particular operator. (Photo by the authors.)
The Data Analyst analyzes the data flowing from the instrument into the data pipeline — usually using some combination of SLAC-developed software (e.g., the AMI, Figure 4, and explained below) and their own experiment-specific software. We call this “Secondary Data Analysis”, and its main goal is to estimate the data quality, that is the signal-to-noise ratio (SNR) coming from the instrument. The data analyst’s job is mostly to calculate this estimate and pass it to the operator and experiment manager.
Figure 4: Examples of many common types of information displays used at LCLS including the AMI and PSANA analysis toolkit.[amifig]
The Experiment Manager uses the information from the data analyst and works with the operator to make decisions about what to do next. The most important decisions (under normal circumstances) about the experiment progress and plan are whether to continue or abort the current measurement, and what sample to measure next. Although each experiment starts out with a specific experiment plan, dynamic (“opportunistic”) re-planning almost always takes place based upon many factors. This will constitute a significant aspect of our model.
Although these three roles are often served by separate people, this is not necessarily so. The roles are often shared in complex ways across fewer or more than three people. For example, the instrument operator may serve as the data analyst, and for in-house experiments the operator might also be the experiment manager. Moreover, with the exception of the instrument operator, the people filling the other two roles may sometimes not even be physically present; They could be involved by phone or screen sharing. Keeping this flexibility in mind, hereinafter we will assume that each role is occupied by a single different individual.
4.3 Typical Experiment Workflow
Figure 5 shows the schedule for LCLS in April 2022. The teams using the beam are color coded by hutch (red is CXI). Experiments with an “X” (as “LX…”) are internal, that is, it is sponsored by SLAC or LCLS, and the team is (usually) composed entirely from SLAC/LCLS staff. Grey blocks represent scheduled downtime, e.g., for maintenance. Note that there are three day shifts and three night shifts. [^8] At the time of this schedule, there were three day and three night shifts. More recently the schedule has been reallocated so that there are only two day and two night shifts. In the present work we are generally focused just on a single operating shift, but it is worth having a sense of the overall structure of experiments.
Figure 5: Example month’s schedule, for April 2022, approximately 27 experiments. [LCLSSchedule]
Once the instrument is prepared and a target is in place the instrument and machine operators begin working together to align the x-ray beam to reach the sample in the desired manner, usually as a stable, focused beam, that data is being taken correctly by the instrument, and that the various downstream data pipelines are working as desired. All of this involves hundreds of physical and computational adjustments. Once alignment is completed the experiment itself can begin, and what follows is the series of measurements that comprise the bulk of the experiment.
5 Cognitive Aspects of LCLS Experiment Operations
LCLS operations are an ideal setting to understand collaborative, real-time, problem-solving. There is an extensive literature in team-based problem-solving that we will not attempt to review here aside from pointing to some recent reviews, e.g., [Graesser2018, Hutchin2010]. Such problem solving is characterized by complex opportunistic interplay of numerous inter-related goals and the physical and cognitive activities that the team deploys to reach these goals.
In acting toward reaching their goals, the team takes advantage of what we will generically call “resources”. The most important category of resources are actions and their affordances (together, “afforded actions”). A slightly too simplistic way to think of what an afforded action is, is as an action that can be accomplished through manipulation of some sort of artifact. The simplest example of this is moving a mirror by virtue of stepping a motor, which is afforded by the motor controls (Figure 3). [^9] There is an important aspect of expertise at play here, which we will not go into much in the present paper. Instead, we assume that the participants in all roles are highly expert in their tasks. This is not the case for what might be called “situation specific” expertise — that is, expertise in how the system happens to be responding, say, today. Collins and Evans have pointed out that expertise is “fractal” [CollinsFractal], but, again, we will not go more deeply into that interesting issue in the present paper. But afforded actions are often much less obvious. For example, particular sorts of displays are artifacts that afford the action of obtaining certain types of information. [^10] We use the term “information resources” throughout to indicate the resources (other than other members of the operating team) that a cognitive agent gathers information from. The most common such resources are visual displays, such as graphs or graphical representation of instruments on computer screens, but there is a wide variety of components from which one gathers information, for example, cameras looking at the instrument itself, auditory signals, paper documents, electronic logs, and even tactile feedback. People can also be thought of, in some cases, providing afforded actions. For example, in the local instrument-side team, the instrument operator’s expertise affords to the experiment manager actions such as getting data and terminating the measurement, and the expertise of the linac operators in the ACR affords various actions to the instrument-side team. Another set of non-obvious actions is afforded by information recorded in the setting through artifacts such as physical and electronic records, such as the eLog, notebook, post-its, and so on utilized around the ICR. These sorts of artifacts afford both information gathering and information storage actions. [^11] Post-Its are quite commonly stuck to various parts of the instrument and controls to facilitate the storage and recovery of important information. The final set of afforded actions that we shall point out are those relating to the problem-solving itself, for example, setting a new goal or declaring a goal to have been reached, rejected, pended, or abandoned. In these cases the “artifact” is a bit more subtle, and might be called the “plan”, although the term “plan” gives this too much of a static feel, whereas in reality the plan is constantly being modified through actions.
Finally, we must point out two important non-physical resources that play a critical role in this setting, as they do in almost all problem-solving settings: time and attention. The operating team must attend to, and respond to, dozens at least, and in some cases hundreds of information resources in order for a measurement to succeed. There are often between 20 and 30 computer windows open across a dozen monitors, and the operator interacts with these partly “demandantly” (that is, responding to status updates and alarms that require a response [Demandance]) and partly intentionally (that is, with the intention of making a change that is not in direct response to, e.g., an alarm). There are many competing goals at play in the positioning of data and control windows, and this often changes during the course of the experiment, as different situations demand attention to, and control inputs to, different aspects of the experiment. The resource of time will play a large role in the present project. However, although attention is extremely important in the fine-grained LCLS operations, we are not going to pay much attention to it in the present paper.
6 LCLS Experimental Operations from the Cognitive Point of View
Next, we describe a simplified timeline of a three-day experiment, and drill down into the measurements, and especially the cognition surrounding measurement-related decision-making. On the first day of the time allocated to a particular experiment, the team gets organized, installing any required special equipment and testing the setup. [^12] Sometimes installation begins before the formally allocated time, if parts of the installation can be done safely while the just previous experiment is still ongoing in a different hutch. The machine operators (in the ACR) get set up to deliver x-rays with the required characteristics. Once that is working, the instrument operators (in the ICR) will test the instrument with unimportant test samples. The team will get their software ready and ensure that it is getting data, and the analysts will conduct analyses on the test samples. Let us assume that all of this takes place on the first day, and that all “real” measurements take place on the second day. This will be described in greater detail below. Throughout these measurements, the collected data is being stored and analyzed at various granularities, also described in more detail below. Finally, on the third and last day of our simplified timeline any specialized equipment added to the instrument will be removed and the remote part of the team departs and the local staff of instrument scientists begin setting up for the next experiment. And so it goes.
In our simplified timeline we have assumed that all the “real” measurements — that is, those that the experimental team will use as actual research data — are carried out on the second day of three in this example. As described briefly above, the “real” samples (as opposed to dummy or control samples) — those from which “real” data is desired — range in importance with respect to the scientific questions at hand. For simplicity let us assume that there are, in this example, 8 “critical” samples, and that 15 minutes of data are intended to be collected on each sample under each of 3 conditions. These details will vary, of course, from experiment to experiment, but a common case is “pumping” the sample with a pulse from a lower energy laser just before the sample is hit by the x-ray beam. This is called the “pump-probe” protocol. The pump — the lower photon energy (e.g., visible) laser pulse — energizes the molecule, changing its confirmation, and then the “probe” from the x-ray laser strikes from the energized molecule, destroying it (because the x-ray energy is so high) but at the same time diffracting and thus providing measurements that can be interpreted later as what amounts to a picture of its energized configuration. [^13] Indeed, this process takes place so rapidly at LCLS that by varying the time-relationship between the pump and probe, one can actually make animations of the conformational changes that molecules undergo during the absorption process. In this sort of experiment, the three desired conditions might be different wavelengths or intensities of the pump pulse, different timings between the pump and the probe, etc.
This whole second day — 24 hours, comprising three 8-hour shifts — might seem like plenty of time to obtain 8x3x15 minutes worth of data, which only comes to a total of 360 minutes, or 6 hours. However, problems inevitably arise. Beam time at LCLS is a very scarce resource, and, as mentioned above, failure to obtain enough good data on the most important samples not only wastes time and money, but could be the difference between a student graduating on time, or even the development of an important drug. At LCLS, problems might arise in any of a number of areas: the Machine itself, which is very complex and dependent on precisely synchronized processes, might be being unstable leading to instability in the x-ray beam delivery; the instrument may be having similar troubles (new parts of it perhaps having just been installed after a cross-country trip for this experiment)141414 This tends to be less of an issue with CXI, x-ray crystallography being a fairly well-worked-out science, although not so much at the energies at which LCLS operates and the pump-probe is quite timing sensitive and can be difficult to get to work well.; the sample may have any of a number of problems, for example, delivery of liquids with various properties tends to result in unpredictable surface tension and turbulence effects. One such problem is central to our model. If enough problems arise, one might run out of time, or even run out of critical samples.
6.1 The Central Decision: What to Measure When and for How Long
In light of these complexities, the decision of what to measure when, and for how long is the most critical decision to be made by the experiment manager. Many factors go into this decision — which is, of course, made in consultation with the rest of the team, and often includes the machine operators in the ACR. To give a sense of the complexities involved, consider that once the team runs out of their allotted beam time (24 hours in this case) they may have to wait months to get additional access if they failed to gather all the desired data. Because of the amount of data involved, and the limited time available to gather it, final (tertiary) data analysis cannot be accomplished in real time. Therefore, the team can only roughly tell, through secondary data analysis, whether the data they are getting is satisfactory. Indeed, this is the specific job of secondary data analysis. For these reasons one would think that it would make the most sense to carry out measurements on the more important sample earlier in the experiment because if problems are leading to lower SNR one can simply take more data.
However, recall that there may be limited sample material, especially of the most important samples, and the problems may resolve, improving the SNR. These considerations vie in the opposite direction, suggesting that if things are not working well (i.e., for whatever reason, the team is seeing low SNR in secondary data analysis), but it is still early in the overall experiment period, it may be better to hold on to the precious critical sample and hope that the problems resolve. But how long do you wait? Half way? Two thirds? Three quarters? And if the data is marginal, and getting worse, should you continue a measurement to wring as much out of the current sample, or simply abort it and move on to more important samples, given an impending hard stop? And once the ongoing measurement is completed, which of the 8x3x15 minute measurements should be tried next? Perhaps even a previous less-than-satisfactory measurement should be repeated?
Obviously, many considerations go into these decisions, and making these decisions as correctly as possible is critical to the success of the experiment. Because of this importance, we choose to focus on this set of decisions in creating our cognitive model of LCLS experiment operations.
7 A Multi-Scale Cognitive Interaction Model of LCLS Instrument Operations
Our goal in the present work is to predict the potential impacts of UI and workflow changes in the most impactful aspects of LCLS operations. Because of the difficulty of carrying out direct user studies, or of employing a “flight simulator” approach, in the LCLS setting, we chose to create a computer simulation (hereafter, “the model”) of some of the most important human-in-the-loop aspects of experiment operations, and in particular of the phenomena relevant to the core decision-making, as discussed above, of when and whether to continue or abort a measurement, and what to measure next. The goal is to use this model to predict the direction and rough shape of changes in experiment efficiency that might result from changes in instrument procedures and/or control interfaces.
In what follows we discuss the central concepts represented in the model. Much more detail is included in the three video supplements to this paper[vids]. The first video is a concept review, and provides more or less the same information as below. The second includes several demonstrations of the model running, and discusses the results depicted in figures 7 through 9. The third video is a deep dive into the model code, which is available on github.[githubmodel].
7.1 Multi-Agent, Multi-Scale
We explicitly model the three primary agents, as described above: The Operator is directly in control of the beamline of the instrument. The Data Analyst observes the AMI and determines when the data is good enough to stop the run as well as gives updates to the Experiment Manager when there is a problem. The Experiment Manager is responsible for the overall control of the experiment and communicates with both the Operator and the Data Analyst to understand the state of the Experiment as well as make changes at a high level to improve the quality and efficiency of the experiment. These roles implicitly create three temporal scales, as depicted in Figure 6.
Figure 6: A diagram of the interaction between the different cognitive scales during the running of the LCLS. This can be understood as Micro is what an individual in the system does and then zoom out for Meso and again for Macro where the logic of the entire experimental task exists. This illustrates how the connection between all of the roles involved influence each other as well as the overall experiment.
Briefly, the three scales of cognitive analysis, depicted in Figure 6, are:
  • Phenomena at the Micro-Cognitive scale are related to detailed controls and displays. Here we see both reasoning and decision-making, as well as skilled performance that we collectively call “peak chasing”. The primary information inputs to peak chasing involve a cognitive parameter called “Functional Acuity” (FA), and parameters relate to the ability to respond to changes efficiently, which we collectively call “Functional Operability” (FO).
  • Phenomena at the Meso-Cognitive scale involve reasoning and decision-making related to the measurement being made now — specifically, whether to continue or abort it.
  • Finally, phenomena at the Macro-Cognitive scale involve reasoning and decision-making related to the course of the experiment, especially regarding how long to gather data on a given sample, and in what order samples should be run.
Although we discuss each of these briefly below, they are described in more detail, and in context, in the supplemental videos[vids].
7.2 The Micro-Cognitive Scale: Peak Chasing, Functional Acuity, and Functional Operability
Real experimental decision-making takes place in a very complex setting, where many extrinsic factors are involved, such as which samples are available, what the value is of those samples, the stability and performance of all aspects of the system, how far through the experiment they are, and how much time the remaining high-value samples are likely to take, as well as large-scale factors such as SLAC and LCLS shift-changes and the requirements for relevant members of the team to take breaks to eat and sleep. Our model, of course, represents a highly simplified version of these, but it does not attempt to take into account all aspects of these factors.
[…]
The operator’s ability to carry out this task is controlled by two parameters: Functional Acuity (FA) and Functional Operability (FO). FA is the operator’s (or, more generally, the team’s) ability to recognize the need to take an action. FO is their ability to take the appropriate action — more precisely, how long it takes to take the appropriate action (and could include terms related to the probability of taking an incorrect action, although we do not model this at present). The term “functional” in these concepts emphasizes that these are not purely cognitive concepts, but are functions of the interaction between individual cognition, team communications, and technical displays and controls. Functional Acuity and Functional Operability have a temporal as well as informational component. For example, in addition to a parameter that controls how far the beam and stream have to be before the operator can notice that an action must be taken (literally the Functional Acuity), there is a Noticing Delay (ND), and then a response delay, and also if the operator has to change which button is being pressed, there is a cost in terms of time based on the distance from the current button to the new one, which is incurred only if one has to switch directions. Again, these are more clearly understandable seeing them in action in the supplemental videos[vids].
Both Functional Acuity and Functional Operability can be thought of as “UI-related” parameters. However, differences in Functional Acuity arise from the interaction between cognition and the display characteristics of the instrument UI, whereas differences in Functional Operability arise from the interaction between cognition and the control characteristics of the instrument UI. Whereas changes to the displays or controls would affect only one or the other of these, more general cognitive interventions, such as alertness or amount of experience with the system will affect both of these parameters.
7.3 The Meso-Cognitive Scale: When to End a Measurement
One of the most important decisions that the operating team must make during an experiment is when to stop taking data on a particular sample, that is, when to end a measurement. In reality, many considerations go into this decision, including the stability and controllability of the x-ray beam, how the data coming from the instruments looks per secondary data analysis combined with the quality of the data that has been obtained so far, additional planning considerations, such as the amount of time remaining in the run and perhaps in the shift, the mental and physical state of the team, how much sample material remains, and considerations relating to the overall experiment (including tertiary data analysis) such as the importance of the sample, and probably other considerations. In our model this decision is made based up the current rate of improvement in SNR, and the beam time remaining.
7.4 The Macro-Cognitive Scale: Sample Choice Planning and Re-Planning
The dynamic environment of an LCLS experiment requires a blend of meticulous pre-planning and continuous real-time, opportunistic decision-making and re-planning. This is markedly different from static, puzzle-like problem-solving where one has the luxury of time and a limited set of variables. The actual measurement process is akin to balancing a pencil on its tip, and is made more complicated by the nature of the novel physics being studied at LCLS, and, of course, by a constant barrage of tiny problems, requiring real-time decision-making and re-planning. While some adjustments, like shortening a measurement if one is obtaining high-quality data, are relatively straightforward, others, such as extending a measurement, or even bringing in additional materials from off-site if necessary, can be more challenging given the strictly limited time frame of LCLS experiments, and are potentially impossible. In our model, the only information used to make these decisions is the information available regarding how well the measurements have been going so far, how much beamtime remains, and the number and quality of samples remaining to test.
7.5 Measurement Error and Performance Quality of Samples
Central to all of the above (and, of course, in the real world LCLS setting) is the quality of the data being gathered. Data quality in settings such as LCLS is often measured in terms of signal-to-noise ratio (SNR), as variously mentioned above. SNR is the slope of a scatter plot of the data, divided by its standard error, and standard error is simply the standard deviation of the data divided by the square root of the number of data points. Therefore, in our model we simply use standard error as a stand-in for (the inverse of) SNR, and the goal of a given measurement is to reach a particular target standard error, which we call “Target Error” (TE), and which is 0.001 by default, for its given sample.
Samples in our model only differ from one another in terms of a single parameter, called “Performance Quality” (PQ). A sample’s PQ is intended to represent knowledge about how easy it is to get data from this sample, based upon either prior experiments, for example in a synchrotron, or theoretical calculation. Higher PQ samples will produce higher quality data, that is, they will reach TE with fewer samples than samples with lower PQ. Somewhat counter-intuitively, more important samples have lower PQ. The reason that lower PQ samples are considered more important is a bit subtle: If it was easy to get data from a given sample — that is, if they have high PQ — then one would have done so already, or one would be sure of their theoretical prediction, and thus it would be less important to get high quality data from samples that have high PQ. An additional complexity arises from the fact that usually the quantity of low PQ (more important) samples on hand is also often limited — again, if samples were unlimited one would have been able to get adequate data on these in a less time-constrained setting, like a ring. One does not want to be wasting rare and important low PQ samples, especially early in the experiment, when the SNR is low for extrinsic reasons, such as experiment stability and operator experience. Similarly, one does not want to be wasting the most important samples (low PQ) stabilizing the instrument. Therefore, one generally will want to run high PQ (less important) samples during stabilization, until one has high confidence in the data coming from the instrument. Put another way, high PQ (less important) samples are those for which you essentially know what data you are likely to get, whereas low PQ (more important) samples are ones that are harder to get data from in either (or both) of theoretical modeling or ring studies. The whole reason for coming to LCLS and running on the XFEL is because it’s hard to get data on the low PQ (more important) samples, otherwise you wouldn’t be at LCLS to begin with.
7.6 Error Adjustment and Cutoff Time
If one lived in a perfect world with unlimited beam time and unlimited samples, any desired specific Target Error (TE) could always be reached. However, in the real world of LCLS, both time and samples are limited commodities. As intimated just above, the team wants to generally run low importance, high PQ, samples early on, during instrument stabilization, and more important, high PQ, samples later, but not so late that they will run out of beam time. In a more realistic model the experiment manager might dynamically make decisions about which sample to measure next, however, our present model is constrained to run samples in High-PQ-to-Low-PQ order. Given this constraint, the only decisions that the experiment manager can make are what TE to shoot for, and when to cut off a given measurement. The model can be set to be constrained regarding the total beam time (called “Cutoff-Time”, where a values of “True” (yes) means that the team will be kicked off the machine after an allotted time), and, similarly, may or may not be allowed to adjust the TE (“Adjust-Error”, where, again, a “True” (yes) values allows the experiment manager to change the TE). The most realistic case, both Cutoff-Time and Adjust-Error will be True (yes), in which case the team may heuristically decide to shoot for a higher or lower TE, depending upon how the measurements are going and the amount of time remaining, but at the end of that time, they will be kicked off the machine.
8 Example Results
In this section we discuss some example results obtained from our model that demonstrate various ways that the mode could be used to investigate changes in the user interfaces or workflows of LCLS instrument-side experiments. These are discussed in more detail in the second of the three auxiliary videos. In each case the vertical axis records the number of data points required to reach the Target Error (default is 0.001), over Performance Quality (PQ on the horizontal axis). In all cases the samples are run from highest to lowest PQ. This is, somewhat confusingly, from the right to the left on all the graphs. Recall that samples with high PQ are less important, and are thus assumed to be used first either as controls, or in order to shake out the instrument. It is actually unrealistic that the most important samples (lowest PQ) would be held until the very end of the experiment, because it is highly likely that the team will run out of beam time and not be able to get enough data for their most important samples. However, allowing the model to re-plan the experiment, resulting in reorganizing the order in which samples are run, would be extremely confusing to try to depict and explain, so we have left this to a separate publication focused just on this topic, where we can explore the effects of sample re-planning in detail.
Figure 7 depicts two explorations with the model. The left graph in Figure 7 scans across three values of Functional Acuity (FA), and is made with Adjust-Error=False and Cutoff-time=False, thus allowing the user to remain as long as necessary to reach the default Target Error (TE) of 0.001. Recall that runs start at high Performance Quality (PQ) and go to progressively lower PQs — that is, from lower importance samples to higher, or right to left on these graphs. Unsurprisingly, it requires less data to reach the TE threshold (0.001) as PQ increases, and as FA improves (becomes smaller), so that the operator notices that the beam and stream are not aligned at smaller distances of misalignment. The right graph in Figure 7 holds FA to 0.1 (the best performing value from the left), and compares Noticing Delays (NDs). In this case, Adjust-Error=True and Cutoff-time=True, which is the most realistic case: The team is going to get kicked off the machine at a fixed time, but they are allowed to adjust the Error Target to try to get some data on all samples. At first glance, in this case, Noticing Delay makes no apparent difference; All the lines are more or less the same. However, as PQ increases, the amount of data gathered appears to be smaller at both high and low PQ, but higher at middle values of PQ. The reason that this happens is that the experiment manager observes that they are going to run out of time with the default TE, and adjusts the TE upward so that some data is obtained for even the last/most important/lowest PQ samples, although they do run out of time in the end anyway. Again, more flexible re-planning would improve the data quality for the most important (lowest PQ) samples, as will be explored in a separate paper.
Figure 7: Left hand: Varying Functional Acuity (FA $\epsilon$ 0.1, 0.5, 1.0) with Adjust-Error=False and Cutoff-time=False. Lower FA corresponds to noticing finer misalignments, and requiring fewer samples to reach the Target Error. Runs start at high Performance Quality (PQ) and go to progressively lower PQs — that is, from lower importance samples to higher, or right to left. With Adjust-Error=False and Cutoff-time=False the experiment is allowed to run to reach TE=0.001 because the team is allowed to continue even after their beam time technically expires. Right hand: Fixed FA = 0.1, varying Noticing Delay (ND $\epsilon$ 1, 5, 10) with ADJUST-ERROR=True and Cutoff-time=True. Higher ND corresponds to slower response to misalignments. ND has much less differential effect at these levels (cf. Fig. 8). With ADJUST-ERROR=True and Cutoff-time=True some data is gathered on all samples, even though the team is kicked off the machine when their beam time expires (cf. Fig. 8 bottom left), although not as much data as if they are allowed to continue, as on the left.
Figure 8 depicts four explorations with the model, scanning Noticing Delay (ND) at three values, and across all combinations of the conditions Adjust-Error (ADJ_) and Cutoff Time (Stop_). The goal of the experimental team is to achieve a good (low) error rate within the allotted time. The Target Error (TE) is the team’s chosen target for a given run. The initial (default) TE is 0.001, and with Adjust-Error(ADJ_)False this never changes. However, with Adjust-Error(ADJ_)True the experiment manager will adjust the TE based on the time remaining and the performance per unit Performance Quality (PQ) to this point. (This is described in detail in the supplemental videos, including giving the algorithm and describing the code.) Recall that in all present simulations we run from higher to lower PQ, that is right to left on these graphs. The first (rightmost) sample error rate is always 0.001. (Note that the data for this rightmost point on each graph is, as expected, almost identical in every case.) When the team is able to adjust the error rate (ADJ_True: top pair [^15] The terms “ADJ” and “Stop” are equivalent to the terms used elsewhere in the paper: “Adjust-error” and “Cutoff-time”, respectively.) the second through last (5th) sample’s TE can be adjusted to try to stay within the given time. When the time is strictly limited Cutoff-time(Stop_)True (left pair in Fig. 7) the team is eventually kicked off the machine. The most realistic condition, is the top left: ADJ_True, Stop_True, and one can see, by comparing the amount of data gathered on the last two sample runs between the top left and bottom left graphs, when the team is allowed to adjust the error target (top left), they do gather some, although not very much, data on the last two (leftmost) critical samples, whereas in they cannot adjust error, but are strictly time-limited (bottom left), essentially no data is gather on the last two (leftmost) critical samples (because they are kicked off the machine!) For comparison, the right pair of graphs are with no time limitation (Stop_False), which is an unrealistic situation, but gives us a sense of how much data needs to be collected to reach a given Target Error. The top right case (ADJ_False, Stop_False) tells us how long it will take to reach 0.001 error, whereas the bottom right case (ADJ_False, Stop_False) gives the times it will take to reach an adjusted TE. These can be compared with their leftward counterparts. Also, the effect of ND is more pronounced in the right hand pair of graphs, because they run significantly longer than those on the left. These results are also discussed in the second half of the “Demo” (2nd) supplemental video[vids].
Figure 8: Three values of Noticing Delay ($\epsilon$ 1, 5, 10) across all combinations of Adjust-Error (ADJ_) and Cutoff Time (Stop_) $\epsilon$ True, False.
Figure 9: The results from the simulation presented in Figure 8 visualize the differences between the conditions of Adjusted Error and Cutoff Time. Left, the simulation executed without any time constraints took considerably longer to yield a result. In contrast, under the time-limited condition, the duration remained constant, irrespective of whether the goal error was modulated or not. Right, the goal error threshold was consistently attained when there was no time restriction, given the unlimited duration available to meet the desired error rate. However, under time constraints, modulating the error resulted in a superior overall goal error in comparison to the non-adjusted error condition. This phenomenon can be attributed to the optimal utilization of time, ensuring that more efficient samples were run while allowing lower goal error for the less efficient ones, thereby enhancing the overall objective error rate.
9 Discussion
Our goal in developing a multi-agent, multi-scale cognitive interaction model of instrument operations at LCLS was to enable the exploration of the impact of proposed UI and workflow/process changes in the human-in-the-loop instrument operations in order to improve the efficiency of results production at LCLS. Scientific user facilities, such as LCLS, are cognitively interesting, and cognitively difficult precisely because they are on the cutting edge of both science and human performance. There are no perfect simulators; If there were one would not need the real machine! And although the instrument scientists are experts at their instruments, they are not experts at each new experiment. Moreover, they are usually working with external users who have widely varying skills, knowledge, and experience, ranging from those who could probably nearly drive the machine themselves, down to grad students, and sometimes even undergrads.
Even though the current model is only a rough analog to the highly complex LCLS setting, it produces sensible results across a variety of parameter ranges that provide insight into how these various factors and scales of cognition interact and impact the efficiency of experimental operations at LCLS.
Modulating Functional Acuity (FA) and Noticing Delay (ND) highlights how improved FA (i.e., faster response to beam misalignment) leads to increased overall experimental efficiency with fewer samples needed for success. In contrast, improved ND produces less significant impacts on efficiency, except in unusual operating regimes of the system, for example when there is no beam time limit. We also investigated quality and duration tradeoffs by adjusting whether the Target Error (our stand in for inverse of SNR) is dynamic or fixed, and by imposing cutoff times, or allowing runs to continue indefinitely. Unsurprisingly, fixed target error requirements produce higher fidelity given unlimited time. However, allowing the target error to adapt yields better time management, ensuring sufficient yields from high performing samples even if lesser quality measurements become truncated by virtue of limited beam time.
These observations showcase how operational efficiency depends substantially on factors related to the human–machine interface, such as Functional Acuity, its operation, such as Functional Operability, and cognitive factors such as attentiveness, amount of experience, and fatigue (represented by Noticing Delay and response delays). Enhancing these aspects of the LCLS setting could optimize scientific productivity.
Previous computational cognitive models have focused on either individual cognition (i.e., how an individual reaches his or her goals, generally through reasoning and action) or on the collaboration between individuals (i.e., how the multiple individuals involved in the activity interact to accomplish the goals). Rarely does a model capture individual cognition at the same time that it captures the closely related cognitive collaboration. Not only does the present model attempt to capture these complexities, but also captures aspects of planning and opportunistic re-planning inherent in complex settings like LCLS. Although only a handful of experiments are presented in the present paper, the model is capable of simulating a much wider scope of parameters, and relatively simple extensions to the existing code can easily broaden its scope even further, even to the level of modeling specific displays, controls, and workflows, and showing how these changes echo at the higher, multi-agent collaboration and experiment planning scales, as is the nature of SLAC/LCLS operations.
Our model merely scratches the surface of the potential of model-based cognitive engineering efforts as applied to scientific user facilities such as LCLS. The present model requires further elaboration and validation with respect to detailed measurements of individual and team behavior. It should be extended to capture more inter- and intra-team communications. Perhaps most importantly, models of learning at all levels should be introduced in order to capture both short term individual learning, for example the instrument scientists becoming familiar with each new experiment; social learning, e.g., the team learning to work together; and longer term learning, such as staff training.
As science probes deeper into the complexities of nature, large scale, beyond-the-state-of-the-art efforts like those at LCLS, and other scientific user facilities, are rapidly increasing in importance. These facilities are in constant flux at every scale. Efforts such as ours that attempt to model the system at multiple scales may be helpful in increasing the efficiency of the science done at these sites by helping to smooth and optimize the human-in-the-loop workflow that will undoubtedly be a significant component of these operations of scientific user facilities for the foreseeable future.
10 Acknowledgements
The research described here arises from a project begun in 2019 by Devangi Vivrekar and Paul Fuoss, with the assistance of Jeff Shrager. The research was picked up by Teddy Rendahl, and then by Wan-Lin Hu and Jonathan Segal. The numerous LCLS and SLAC staff and visiting scientists that permitted us to observe them working, and sat for interviews over those years, are too numerous to mention, but we are especially grateful to Andy Aquila, Mark Hunter, Meng Liang who welcomed us into their experiments, and to Dan Flath and especially Jana Thayer for valuable discussions and support. We also thank Nick Briggs for a careful reading that uncovered numerous typos. This project was supported by the U.S. Department of Energy, Office of Science, Office of Basic Energy Sciences under Contract No. DE-AC02-76SF00515.
11 Contributions of Authors
Jonathan Segal did most of the model coding, ran the simulations, and created the graphics. Wan-Lin Hu did most of the data collection and analysis. Frank Ritter provided detailed guidance on applied cognition and human factors. Paul Fuoss conceived the project, obtained the funding, dealt with all administrative aspects, and was the primary applied physics informant. Jeff Shrager managed project, participated in the observations and model coding, and wrote this paper. All authors reviewed and commented on the paper.
[…]